OpenStudy (mathhelp346):

A design calls for a rectangular garden with an area of 180 square feet. If the length is to be 3 feet more than the width, set up an equation and solve that equation to find the dimensions of the garden.

OpenStudy (mathhelp346):

I got w=l-177 and l=w+177, but that doesn't sound right

OpenStudy (mathhelp346):

why is that one w+3? shouldn't it be l+3?

OpenStudy (anonymous):

 length is w + 3